# A Systematic Investigation of Knowledge Retrieval and Selection for Retrieval Augmented Generation

### v1

@@ 6 Conclusion and Discussion
In this study, we systematically examined the behavior of RAG generation in relation to the quality of retrieved knowledge.
@@ 6.1 Overall Observations
Summing up the results in Section 5, we find an interaction effect between generator capability and the ambiguity of the task/dataset on the RAG generation performance. We find two types of generator behavior:
A strong generator model can achieve good performance without a knowledge selector because it is robust to noise, and thus performs better the more gold knowledge is retrieved, despite the presence of distractor knowledge. Thus, knowledge recall correlates well with generation F1, and there is not much room for improvement through the addition of a knowledge selector.
@@ 6.2 Recommendations to Future Practitioners
Based on our observations, we make the following recommendations for future practitioners considering how to improve performance in a real-world RAG scenario.
First, benchmark the generation performance without external knowledge, with all candidate knowledge, and if possible, with gold knowledge only. The “no knowledge” setting measures the generator’s base performance, “full knowledge” serves as a strong baseline corresponding to knowledge retrieval with perfect recall, and “gold knowledge” gives the upper bound of RAG performance with perfect knowledge selection. The gap between “full knowledge" and “gold knowledge" is the potential performance gain brought by adding a knowledge selector.
To improve RAG performance, increasing the knowledge recall score is the most effective strategy. Thus, for modern LLM generators, improving the knowledge retriever’s recall is key. In practice, the retriever or selector may encounter false-negative gold knowledge during training, as is the case with WoW, which makes prioritizing recall even more important. Moreover, because modern LLM generators have long maximum input windows (e.g. 128k for GPT-4o & -mini), having too much knowledge is less likely to be a problem.
Only when the knowledge recall score is high can a knowledge selector, which increases knowledge precision, be potentially helpful. Moreover, if the recall is too low, the downstream generation performance may not see any benefits from a knowledge selector with middling performance — it may even be harmed; only when the knowledge selector’s precision is very high will the overall performance improve.
@@
In addition to the cost issue, there are very few RAG datasets with human-annotated knowledge Friel et al. (2024), which further limits our simulation experiment settings. Furthermore, even though we regard WoW as a relatively noisy dataset in this work compared to HotpotQA, to the best of our knowledge, WoW is one of the most cleanly annotated datasets among all datasets. Therefore, given our hypothesis in Section 5 that WoW has a larger solution space than HotpotQA, we do not have numerical evidence because each example in WoW only contains one gold response thus we cannot measure the size of the solution space.
@@ Appendix A Appendix
A.1 Prompts
@@
Tables 3 & 4 show the prompts we use for response generation on WoW and HotpotQA, respectively. Tables 5 & 6 show the prompts we use for the “LLM selector" knowledge selection baseline on WoW and HotpotQA, respectively. While we experimented with Chain-of-Thought Wei et al. (2022) for these prompts, they did not outperform zero-shot prompting for LLaMA 3.1 8B and Mistral-7B-Instruct, so we use zero-shot prompting throughout our experiments to keep the settings as simple as possible.
A.2 Additional Plots
@@
Figures 7, 8, and 9 extend Figure 3, with knowledge precision, knowledge recall, and knowledge F1 plotted against response/answer F1 for WoW and HotpotQA, respectively, for all three LLM generators. Figure 8 shows that knowledge recall correlates strongly with response/answer F1 for both GPT-4o-mini and LLaMA 3.1 8B on both WoW and HotpotQA, as well as for Mistral 7B-Instruct on WoW, while Figure 9 shows that knowledge F1 has the stronger correlation for Mistral 7B-Instruct on HotpotQA.
A.3 Noisy HotpotQA
@@
As we discuss in Section 4.1, the WoW gold knowledge annotations are noisy in that only one sentence can be labeled as gold knowledge, even if other sentences are still relevant. To test the impact of such knowledge annotation noise on our RAG simulation results, we deliberately inject noise into HotpotQA: we drop one sentence from the gold knowledge sentences for each example to create a noisy HotpotQA dataset. As a result, the average number of gold sentences per example drops from 2.4 to 1.4, and a small portion of the “distractor" knowledge in the noisy dataset is actually gold knowledge from the original HotpotQA (i.e. false-negative knowledge).
@@
As Figures 5 & 6 show, the injected noise drastically impairs the knowledge selector’s efficacy: even given a “perfect” knowledge selector with a knowledge precision score of 100%, the answer F1 still underperforms the “full knowledge” setting. This is because our knowledge precision and recall scores are computed based on the labeled gold knowledge sentences, so if the labels are noisy, the knowledge metrics are corrupted while the answer F1 is intact. Note that Figure 6 presents a non-monotonic trend in improving knowledge selector, just as we observed in Section 5 for WoW, verifying that noisy knowledge annotations cause this non-monotonic behavior.
@@
Prompt
The conversation is about “{{persona}}”.
{if history %}
Here is the conversation history:
{for turn in history %}
{{turn.speaker}}: {{turn.text}}
{endfor %}
{endif %}
{if context %}
Title: {{ context.title }}
Sentences:
{for index, sentence in context.sentences | enumerate %}
- {{ index + 1}}. {{ sentence }}
{endfor %}
Answer the decision of selecting each sentence with a list of boolean values. Make sure the length of the list is the same as the number of sentences.
Answer in JSON format so that I can use Python eval() to parse it, e.g.,:
“answer”: [True, False, …, True, False],
Your answer:
Table 5: Jinja2 prompt template for Wizard of Wikipedia knowledge selector.
Prompt
Given this question from HotpotQA:
{{ question }}
For each of the following sentences in the context, judge whether they are relevant to answering the question.
Title: {{ context.title }}
Sentences:
{for index, sentence in context.sentences | enumerate %}
- {{ index + 1}}. {{ sentence }}
{endfor %}
Answer the decision of selecting each sentence with a list of boolean values. Make sure the length of the list is the same as the number of sentences.
Answer in JSON format so that I can use Python eval() to parse it, e.g.,:
“answer”: [True, False, …, True, False],
Your answer:
Table 6: Jinja2 prompt template for HotpotQA knowledge selector.
Figure 7: Scatter plot of answer F1 versus the knowledge precision for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA.
Figure 8: Scatter plot of answer F1 versus the knowledge recall for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA.
Figure 9: Scatter plot of answer F1 versus the knowledge F1 for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA.

### v2

@@ 6 Conclusion and Discussion
In this study, we systematically examined the behavior of RAG generation in relation to the performance of knowledge retrieval and selection.
@@ 6.1 Overall Observations
Summing up Section 5, we find an interaction effect between generator capability and ambiguity of the task/dataset on RAG generation performance. We find two types of generator behavior:
A strong generator model can achieve good performance without a knowledge selector because it is robust to noise, and therefore performs better as more gold knowledge is retrieved, despite the presence of distractor knowledge. Thus, knowledge recall correlates well with generation F1, and there is less room for improvement through the addition of a knowledge selector.
@@ 6.2 Recommendations to Future Practitioners
Based on our observations, we make the following recommendations for future practitioners considering using knowledge selectors to improve performance in a real-world RAG scenario:
Benchmark the generation performance without external knowledge, with all candidate knowledge, and if possible, with gold knowledge only. The “no knowledge” setting measures the generator’s base performance, “full knowledge” serves as a strong baseline corresponding to knowledge retrieval with perfect recall, and “gold knowledge” gives the upper bound of RAG performance with perfect knowledge selection. The gap between “full knowledge" and “gold knowledge" is the potential performance gain brought by adding a knowledge selector.
To improve RAG performance, increasing the knowledge recall score is the most effective strategy. Thus, for modern LLM generators, improving the knowledge retriever’s recall is key. In practice, the retriever or selector may encounter false-negative gold knowledge during training, as is the case with WoW, which makes prioritizing recall even more important. Moreover, because modern LLM generators have long maximum input windows (e.g. 128k for GPT-4o & -mini), having too much knowledge is less likely to be a problem. If the number of knowledge sentences must be constrained to reduce computational cost, the top-$k$ knowledge sentences should maintain the same level of knowledge recall and precision as the non-length-constrained settings.
Only when the knowledge recall score is high can a knowledge selector, which increases knowledge precision, be potentially helpful***In a length-constrained scenario, if we consider $k$ randomly sampled knowledge sentences as a baseline, rather than the “full knowledge” setting, then the knowledge selector is more likely to be helpful due to weaker baseline performance.. Moreover, if the recall is too low, the downstream generation performance may not see any benefits from a knowledge selector with middling performance — it may even be harmed; only when the knowledge selector’s precision is very high will the overall performance improve.
@@
In addition to the cost issue, there are very few RAG datasets with human-annotated knowledge Friel et al. (2024), which further limits our simulation experiment settings. Furthermore, even though we regard WoW as a relatively noisy dataset in this work compared to HotpotQA, to the best of our knowledge, WoW is one of the most cleanly annotated datasets among all datasets. We cannot verify our hypothesis in Section 5 that WoW has a larger solution space than HotpotQA without re-annotating the dataset because each example in WoW only contains one gold response. Our experiments in Appendix A.4 compare the noisiness of gold knowledge annotations in WoW with those in HotpotQA.
Uniform Sampling Probably for Simulation.
In Section 3.2, we draw gold and distractor knowledge from a uniform distribution $p_{gold}$ and $p_{noise}$ for simplicity since we do not have a prior assumption of the knowledge selector’s preference. However, a real knowledge selector may be more likely to select one knowledge sentence than another. Nonetheless, we successfully identified the knowledge precision and recall scores together as good predictors of the generation performance as we analyzed in Section 5.
@@ Appendix A Appendix
A.1 Implementation Details
We do not perform any fine-tuning or hyperparameter tuning. We set the LLMs’ temperature to 0 and ensure that our zero-shot generation prompts (Appendix A.2) are shorter than their maximum input lengths. We use the API services of OpenAI gpt-4o-mini-2024-07-18***https://platform.openai.com/docs/models/gpt-4o-mini, Together AI***https://docs.together.ai/docs/chat-models meta-llama/Meta-Llama-3.1-8B-Instruct-Turbo, and mistralai/Mistral-7B-Instruct-v0.1. We report results on the first 500 examples in the HotpotQA training set and the first 100 conversations (452 wizard utterances) from the “test seen” set of WoW. Our set of experiments cost about 50 USD from OpenAI and about 50 USD from Together.ai.
A.2 Prompts
@@
Tables 3 & 4 show the prompts we use for response generation on WoW and HotpotQA, respectively. While we experimented with Chain-of-Thought Wei et al. (2022) for these prompts, they did not outperform zero-shot prompting for LLaMA 3.1 8B and Mistral-7B-Instruct, so we use zero-shot prompting throughout our experiments to keep the settings as simple as possible.
A.3 Additional Plots
@@
Figures 10, 11, and 12 extend Figure 3, with knowledge precision, knowledge recall, and knowledge F1 plotted against response/answer F1 for WoW and HotpotQA, respectively, for all three LLM generators. Figure 11 shows that knowledge recall correlates strongly with response/answer F1 for both GPT-4o-mini and LLaMA 3.1 8B on both WoW and HotpotQA, as well as for Mistral 7B-Instruct on WoW, while Figure 12 shows that knowledge F1 has the stronger correlation for Mistral 7B-Instruct on HotpotQA.
A.4 Noisiness of Wizard of Wikipedia vs. HotpotQA
To intuitively compare the noisiness of WoW vs. HotpotQA, for each example, we feed each individual candidate knowledge sentence, regardless of gold or distractor status, to the GPT-4o-mini generator model and measure the answer F1 score to measure how each individual knowledge sentence affects the generation performance. As Figure 7 shows, WoW knowledge sentences result in a wide, continuous distribution of response F1 scores, indicating that even the “distractor" knowledge positively contributes to the gold response to some extent. In contrast, we find that 70% of HotpotQA sentences are true distractors that cause the generator to produce incorrect answers with 0 F1 score, while 20% of sentences result in correct answers. These results indicate that WoW is a much noisier dataset than HotpotQA.
A.5 Noisy HotpotQA
@@
As we discuss in Section 4.1, the WoW gold knowledge annotations are noisy in that only one sentence can be seelcted as gold knowledge, even if the other retrieved sentences are still relevant. To test the impact of such knowledge annotation noise on our RAG simulation results, we deliberately inject noise into HotpotQA by removing one sentence from the gold knowledge set for each example to create a noisy HotpotQA dataset. As a result, the average number of gold knowledge sentences per example drops from 2.4 to 1.4, and a small portion of the “distractor" knowledge in the noisy dataset is actually gold knowledge from the original HotpotQA (i.e. false-negative knowledge).
@@
As Figures 5 & 6 show, the injected noise drastically impairs the knowledge selector’s efficacy: even given a “perfect” knowledge selector with a knowledge precision score of 100%, the answer F1 still underperforms the “full knowledge” setting. This is because our knowledge precision and recall scores are computed based on the labeled gold knowledge sentences, so if the labels are noisy, the knowledge metrics are corrupted, while the answer F1 is intact. Note that Figure 6 presents a non-monotonic trend in improving knowledge selector, just as we observed in Section 5 for WoW, verifying that noisy knowledge annotations cause this non-monotonic behavior.
@@
A.6 Length-Constrained Knowledge Selection and Answer Generation
To study the effect of length-constrained knowledge inputs on answer generation performance, we limit the number of knowledge sentences to $k$ by randomly sub-sampling $k$ sentences when more than $k$ sentences are selected by the simulated knowledge selector***This random sampling does not break the experiment because our knowledge retriever and selector are simulated via random sampling, which is not the case for actual model-based retrievers and selectors.. We choose $k=3$ for both WoW and HotpotQA because their average number of candidate knowledge sentences before random sampling are 7.1 and 9.5 respectively.
As Figure 8 shows, the length of the knowledge input mostly correlates with knowledge precision. As Figure 9 shows, constraining the knowledge input size pushes the points in the upper-left corner of the plots in Figure 2 into the rest of the space while keeping the overall trend of the points the same. Our observations in Section 5 hold even with length-constrained knowledge selection: a sufficiently high knowledge recall score is a pre-requisite for a knowledge selector outperform the “full knowledge” setting.
@@ A.6 Length-Constrained Knowledge Selection and Answer Generation
Figure 7: Histogram of answer F1 distributions of Wizard of Wikipedia and HotpotQA by feeding each individual candidate sentence to the GPT-4o-mini generator.
Figure 8: Scatter plot of the number of selected knowledge sentences versus the knowledge precision on Wizard of Wikipedia (left) and HotpotQA (right). Each data point corresponds to a full experiment on the entire sampled dataset.
Figure 9: Scatter plot of answer F1 versus the knowledge precision for GPT-4o-mini on Wizard of Wikipedia (left) and HotpotQA (right) by limiting the number of knowledge sentences up to k=3. The dots highlighted in orange indicate settings outperforming the “full knowledge” setting (without constraining the number of sentences for fair comparison with other figures), while those highlighted in cyan indicate settings underperforming the “no knowledge" setting. Each data point corresponds to a full experiment on the entire sampled dataset.
@@
Figure 10: Scatter plot of answer F1 versus the knowledge precision for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA. Each data point corresponds to a full experiment on the entire sampled dataset.
Figure 11: Scatter plot of answer F1 versus the knowledge recall for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA. Each data point corresponds to a full experiment on the entire sampled dataset.
Figure 12: Scatter plot of answer F1 versus the knowledge F1 for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA. Each data point corresponds to a full experiment on the entire sampled dataset.